# tableReferences for Table S1

### v1

@@ 2.2 Classifier-based: Classifier Two-Sample Test (C2ST)
The Classifier Two-Sample Test (C2ST) uses a classifier that discriminates between samples from two distributions (Fig. 4a) (Lopez-Paz & Oquab, 2016, Friedman, 2003). The distance between the distributions can then be quantified with various measures of classifier performance. The classification accuracy provides a particularly intuitive and interpretable measure of the similarity of the distributions. If the classification accuracy is 0.5, i.e., the classifier is at chance level, the distributions are indistinguishable to the classifier (Fig. 4c, left), while higher accuracy indicate differences in the distributions (Fig. 4c, middle). If the C2ST is 1.0, the two distributions have no (or very little) overlap in their supports (Fig. 4c, right). Given two distributions, the C2ST has a ‘true’ (optimal) value, which is the maximum classification accuracy attainable by any classifier (Fig. 4a). This optimal value can be computed if both distributions allow evaluating their densities, but this is not usually possible if only data samples are available. In that case, one aims to train a classifier, such as a neural network (Fig. 4b), that is as close to the optimal classifier as possible.
One of the main benefits of the C2ST is that its value is highly interpretable (the accuracy of the classifier). C2ST can also be used to test the statistical significance of the difference between two sets of samples. Unlike other measures, however, C2ST can be expensive to compute because it requires training a classifier and using the classification accuracy as a differentiable training objective is not straightforward (see Section A.5.1). Furthermore, the value is dependent on the capacity of the classifier, and hence on many hyperparameters such as classifier architecture or training procedure. In our experiments, we used a scikit-learn Multi-Layer-Perceptron classifier, combined with a five-fold cross-validation routine to estimate the accuracy returned (Pedregosa et al., 2011).
@@
As mentioned above, for any realistic scenario, the C2ST is computed by training a classifier. The resulting C2ST will only be a good measure of distance between real and generated data if the classifier is close or equal to the optimal classifier. To demonstrate the behavior of the C2ST if this is not the case, we fitted a Gaussian distribution to data that was sampled from a Mixture of Gaussians (Fig. 5a, left). The optimal C2ST between these two distributions is 0.65 (which can be computed because Gaussians and Mixtures of Gaussians allow evaluating densities). If the C2ST is estimated with a neural network, however, we observe that this C2ST can be systematically underestimated: for example, when only few samples from the data and generative model are available, the neural network predicts a C2ST of 0.5 (Fig. 5a, top right) – in other words, it predicts that the generative model and the data follow the same distribution. Similarly, if the neural network is not expressive enough, e.g. with too few hidden units, the classifier will return a low C2ST, around 0.5 (Fig. 5a, bottom right). These issues can make the C2ST easy to misuse: In many cases, reporting a low C2ST is desirable for generative models since it indicates that the model perfectly matches data, but one can achieve a low C2ST simply by not investing sufficient time into obtaining a strong classifier.
@@
We previously argued that the C2ST is an interpretable measure – while this is generally true, the C2ST can sometimes be surprisingly high even if the generative model seems well aligned with the data. For example, when the generative model aligns very well with the data for every marginal, the C2ST can still be high if the data is high-dimensional (Fig. 5b). Because of this, it can be difficult to achieve low C2ST values on high-dimensional data. To further demonstrate this, we fitted a Gaussian distribution and a mixture of 20 Gaussian distributions to the ‘ones’ of the MNIST dataset. Although the Mixture of Gaussians (Fig. 5c, bottom row) looks better than a single Gaussian (Fig. 5c, middle row), both densities have a C2ST of 1.0 to the data (obtained with a ResNet on $\approx$4k held-out test datapoints).
@@
While we focus on a standard C2ST definition by using classification accuracy as the C2ST distance (Lopez-Paz & Oquab, 2016), any other performance metric for binary classification could be used (Raschka, 2014). Kim et al. (2019) even argue that classic accuracy is sub-optimal due to the “binarization” of the class probabilities and proceeds to instead use the mean squared error between the predicted and ‘target’ value of 0.5. Other approaches instead construct a likelihood ratio statistic (Pandeva et al., 2022). Additionally, instead of using the estimated class probabilities, Cheng & Cloninger (2022) consider using the average difference in logits (i.e., activations in the last hidden layer). However, classification accuracy is still the most commonly used variant of C2ST.
@@ 2.3 Kernel-based: maximum mean discrepancy (MMD)
Figure 6: Maximum mean discrepancy (MMD). (a) Two example distributions $p_{1}(x)$, $p_{2}(x)$ and observed data $p_{true}(x)$ that we want to compare. (b) MMD can be defined as the difference between the expectations of some embedding function $\phi(x)$. If we take the identity as embedding ($\phi^{(1)}(x)=x$; left), we end up computing the differences between the means of the distributions, which are all equal for the three distributions. If we add a quadratic feature ($\phi(x)^{(2)}=[x,x^{2}]^{\mathsf{T}}$; right), we can distinguish distributions with different variances. Note that we still have $\mathsf{MMD}^{2}[\phi^{(2)},p_{2},p_{true}]=0$, despite $p_{2}$ being different from $p_{true}$ (c) Using the kernel trick we can avoid computing the embeddings all together but use implicit embeddings that capture all relevant features of the distributions.
MMD is a popular distance metric that is applicable to a variety of data domains, including high-dimensional continuous data spaces, strings of text as well as graphs (Borgwardt et al., 2006, Gretton et al., 2012a, Muandet et al., 2017). It has been used to evaluate generative models (Sutherland et al., 2021, Borji, 2019, Lueckmann et al., 2021) and also has the ability to indicate where the model and the true distribution differ (Lloyd & Ghahramani, 2015). The distance provided by MMD can straightforwardly be used to test whether the difference between two sets of high-dimensional samples is statistically significant (Gretton et al., 2012a).
To assess whether two set of samples are drawn from the same distribution, MMD makes use of a kernel function to (implicitly) embed the samples via an embedding function $\phi$, also called feature map. If we choose the right kernel, we can end up embedding our samples in a space where properties of the underlying distributions are easily compared. We will motivate the use of the kernel in MMD by illustrating different explicit embeddings before introducing the implicit embedding via a kernel $k$. Note that this explanation is inspired by Sutherland (2019).
@@
If we want to compare samples of real numbers from two distributions $p_{1}$ and $p_{2}$ (Fig. 6a), we can think about different embedding functions $\phi$ to compare these. The simplest possible function $\phi^{(1)}:\mathbb{R}\rightarrow\mathbb{R}$ is the identity mapping $\phi^{(1)}(x)=x$ (Fig 6b, left). However, in this case the MMD will simply be the absolute difference between the means (first moments) of the distributions (for details, see Section A.2):
@@
As both models and the true distribution in Fig. 6 have the same mean, this does not yet let us discriminate between them. If we now expand our embedding with a quadratic term, $\phi^{(2)}:\mathbb{R}\rightarrow\mathbb{R}^{2}$ as $\phi^{(2)}(x)=\begin{bmatrix}x\\ x^{2}\end{bmatrix}$ (Fig 6b, right), the MMD yields (for details, see Section A.2)
@@
We can now rewrite the inner product $\langle\phi(x),\phi(x^{\prime})\rangle$ in terms of a kernel function $k$: $\langle\phi(x),\phi(x^{\prime})\rangle=k(x,x^{\prime})$. Thus, if we can find a kernel for our feature map, we can avoid explicitly computing the features altogether but instead, we directly compute
@@
Evaluating the kernel function instead of explicitly calculating the features is often called the kernel trick (Fig. 6c). If we can define a kernel whose corresponding embedding captures all, potentially infinitely many moments, we would have an MMD that is zero only if two distributions are exactly equal. These kernels are called characteristic (Section A.2,  Gretton et al. 2012a), and include the commonly used Gaussian kernel: $k_{G}(x,x^{\prime})=\exp(-\frac{\lVert x-x^{\prime}\rVert^{2}}{2\sigma^{2}})$ (see, e.g., Sriperumbudur et al. (2009) for other characteristic kernel choices).
@@
Typically the kernel version of MMD is used, which is straightforwardly estimated with its empirical (unbiased) estimate:
@@
MMD in this form can be applied to many forms of data, as long as we can define a kernel, which can include graphs (Vishwanathan et al., 2010, Gärtner, 2003) or strings of text (Lodhi et al., 2002), in addition to vectors and matrices.
When we estimate the MMD with a finite number of samples, the selection of the right kernel and its parameters becomes crucial. For example, when using a Gaussian kernel, one has to choose the bandwidth $\sigma$. The MMD approaches zero if we take $\sigma$ to be close to zero (then $k_{G}(x,x^{\prime})=1\;\mathsf{if}\;x=x^{\prime}\;\mathsf{else}\;k_{G}(x,x^{\prime})\rightarrow 0$) or if $\sigma$ is large (then $k_{G}(x,x^{\prime})\rightarrow 1\;\forall\;x,x^{\prime}$(Gretton et al., 2012a). A common heuristic to remedy this parameter choice is picking the bandwidth based on the scale of the data. The median heuristic set the bandwidth to the median distance between points in the aggregate sample (Gretton et al., 2012a). Another common approach is based on cross-validation, or data splitting (Gretton et al., 2012a; b, Jitkrittum et al., 2016, Sutherland et al., 2021): The dataset is divided, with a hold-out set used for kernel selection, and the other part used for evaluating MMD. While the data splitting method does not involve any heuristic, it can lead to errors in MMD since it reduces the number of data points available for estimating the MMD. Recent work attempts to choose hyperparameters without employing data splitting or any heuristic (Biggs et al., 2023, Schrab et al., 2023, Kübler et al., 2022b; a).
While we aim in general for a kernel that captures the (dis)similarity between the data points well, such a kernel can be domain specific or specifically designed for downstream analysis tasks. The similarity between two strings (e.g., DNA sequences, text) can for instance be estimated by looking at the frequency of small subsequences (Leslie et al., 2001, Lodhi et al., 2002). It is furthermore possible to aggregate simpler kernels into a more expressive one (Gretton et al., 2012b), or to use a deep kernel (i.e., based on neural networks) that can exploit features of particular data modality such as images (Liu et al., 2020, Gao et al., 2021).
@@ 2.4 Network-based: Embedding-space measures
Distribution comparisons on structured data spaces, such as the set of natural images, present unique challenges. Such data is usually high-dimensional (high-resolution images) and contains localized correlations. Furthermore, images of different object classes (such as airplanes and dogs) share low-level features in the form of edges and textural details but differ in semantic meaning. Similar challenges occur for time-series data, natural language text, and other complex data type (Smith & Smith, 2020, Jeha et al., 2021).
@@
Figure 7: Network-based metrics. Instead of directly computing distances in data space, complex data e.g., natural images of dogs sampled from $p_{1}(x)$ and aircraft sampled from $p_{2}(y)$, are jointly embedded into a vector space. The embedding function can, for example, be a deep neural network. The resulting distributions in feature space are then compared by a classical measure of choice $D$.
@@
In this section, we take the example of natural images, but the presented framework generalizes to other data types. Naive distances would operate on a per-pixel basis, leading to scenarios where, for example, white dogs and black dogs are considered vastly different despite both being categorized as dogs. As we would like to have a distance measure that operates based on details relevant to the comparison we can leverage neural networks trained on a large image dataset that captures features ranging from low-level to high-level semantic details: While earlier layers in a convolutional neural network focus on edge detection, color comparison, and texture detection, later layers learn to detect high-level features, such as a dog’s nose or the wing of an airplane, which thought to be relevant for a meaningful comparison. Embedding-based distances use these activations of neural network layers as an embedding to compare the image distributions. The most popular distance in this class is the Fréchet Inception Distance (FID) (Heusel et al., 2017), used to evaluate generative models for images. The FID uses a convolutional neural net’s embeddings (specifically InceptionV3 (Szegedy et al., 2015a)) to extract the relevant features, applies a Gaussian approximation in the embedding space, and computes the Wasserstein distance on this approximation.
@@
An FID-like measure, in essence, requires a suitable embedding network $f:\mathcal{X}\rightarrow\mathbb{R}^{d}$, where $f$ transforms the data from the original high-dimensional space $\mathcal{X}$ into a lower-dimensional, feature-rich representation in $\mathbb{R}^{d}$ (Fig. 7). Once the data samples are mapped into this reduced space through the embedding network, the two sets of embedded samples can be compared using the appropriate distance. When evaluating generative models for natural images, it is common to approximate the embedded distributions with Gaussian distributions by estimating their respective mean $\mu$ and covariances $\Sigma$. Under this Gaussian approximation, the squared Wasserstein distance (also known as the Fréchet distance) can be analytically computed as
$W^{2}((\mu_{1},\Sigma_{1}),(\mu_{2},\Sigma_{2}))=\|\mu_{1}-\mu_{2}\|^{2}+\text{Tr}\left(\Sigma_{1}+\Sigma_{2}-2\left(\Sigma_{1}\Sigma_{2}\right)^{\frac{1}{2}}\right).$
@@
In principle, any appropriate metric can be used in place of the Fréchet distance. Jayasumana et al. (2023) show that the MMD (Section 2.3) can be better suited as a metric in the embedding space, as it perceptually matches human judgement on assessing image quality and coverage in generative models. A related and commonly used quality measure for images is the Inception Score (Salimans et al., 2016). In contrast to the FID, this measure uses the average InceptionV3 predicted class probabilities and compares them with the true marginal class distribution. Note that while both this score and the FID can agree with traditional distances (e.g. certain divergences), they might evaluate models differently (Betzalel et al., 2022); see Barratt & Sharma (2018) for further limitations of the Inception Score.
@@
One of the biggest limitations is the requirement of a suitable embedding net. Newer and more robust networks, such as the image network of the CLIP (Radford et al., 2021) vision-language model, provide better and semantically more consistent embeddings (Betzalel et al., 2022) than the InceptionV3 network. However, as the embedding network is generally non-injective, identical distributions in the embedding space may not necessarily translate to identical distributions in the original space. Previous research has demonstrated the FID’s sensitivity to preprocessing such as image resizing and compression (Parmar et al., 2022). Additionally, FID estimates are biased for finite sample sizes, making comparisons unreliable due to dependency on the generative model. However, methods to obtain a more unbiased estimate have been proposed ($\mathsf{FID}_{\infty}$; Chong & Forsyth 2020, Betzalel et al. 2022).
3 Comparison and scalability
When evaluating (or training) generative models, it is important to understand that different statistical distances penalize different features of the generated samples. They might for instance weigh differently how important it is to have large sample variability versus how well the modes of the true distribution are captured (as can be illustrated by optimizing a mis-specified model using the different distances, see Fig. S1 and  Theis et al. 2016). Thus, using different statistical distances to evaluate the quality of generative models can lead to different conclusions.
Figure 8: Scalability of different statistical distances with sample size and dimensionality. (a,b) Comparison of sample sets with varying sample size (between 50 and 4k samples per set) of a ‘true’ distribution either with a second dataset of the same distribution or with a sample set from an approximated/shifted distribution. We show the mean and standard deviation over five runs of randomly sampled data. Note that the subscript for MMD distances (bottom) denotes the bandwidth of the Gaussian kernel used for a given dataset and we report the squared distance for MMD. (a) Distances for the 2d-MoG example shown in Fig. 1 compared to samples from a unimodal Gaussian approximation with the same mean and covariance. (b) Distances for a ten-dimensional standard normal distribution, for which the first dimension is shifted by one for the shifted dataset. (c) Distances based on 10k samples from a standard normal distribution with varying dimensions (between 5 and 1000). As in (b), the first dimension is shifted by one for the ‘shifted’ dataset. We show the mean and standard deviation over five runs of randomly sampled data. One MMD bandwidth was selected for all n-dimensional datasets.

### v2

@@ 2.2 Classifier-based: Classifier Two-Sample Test (C2ST)
The Classifier Two-Sample Test (C2ST) uses a classifier that discriminates between samples from two distributions (Fig. 4a;  Lopez-Paz & Oquab 2017, Friedman 2003). The distance between the distributions can then be quantified by various measures of classifier performance. For example, one would train a classifier $c(x)$ to distinguish samples from the generative model and the data, and then evaluate the C2ST as $\frac{1}{2}[\mathbb{E}_{p(x)}[\mathds{1}(c(x)=0)]+\mathbb{E}_{q(x)}[\mathds{1}(c(x)=1)]]$. The classification accuracy provides a particularly intuitive and interpretable measure of the similarity of the distributions. If the classification accuracy is 0.5, i.e., the classifier is at chance level, the distributions are indistinguishable to the classifier (Fig. 4c, left), while higher accuracy indicate differences in the distributions (Fig. 4c, middle). If the C2ST is 1.0, the two distributions have no (or very little) overlap in their supports (Fig. 4c, right). Given two distributions, the C2ST has a ‘true’ (optimal) value, which is the maximum classification accuracy attainable by any classifier (Fig. 4a). This optimal value can be computed if both distributions allow evaluating their densities, but this is not usually possible if only data samples are available. In that case, one aims to train a classifier, such as a neural network (Fig. 4b), that is as close to the optimal classifier as possible.
One of the main benefits of the C2ST is that its value is highly interpretable (the accuracy of the classifier). C2ST can also be used to test the statistical significance of the difference between two sets of samples. However, calculating C2ST can be resource-intensive because it involves training a classifier. Additionally, using classifiers in a differentiable training objective can be challenging (but is sometimes possible, as in, e.g., Generative Adversarial Networks;  Goodfellow et al. 2014). Furthermore, the value is dependent on the capacity of the classifier, and hence on many hyperparameters such as classifier architecture or training procedure. This dependence on a trained classifier can result in C2ST estimates that are biased, and the variety of possible classifier architectures means theoretical guarantees such as sample complexity are difficult to determine. In our experiments, we used a scikit-learn Multi-Layer Perceptron classifier, combined with a five-fold cross-validation routine to estimate the accuracy returned (Pedregosa et al., 2011).
@@
As mentioned above, for any realistic scenario the C2ST is computed by training a classifier. The resulting C2ST will only be a good measure of distance between real and generated data if the classifier is close or equal to the optimal classifier. To demonstrate the behavior of the C2ST if this is not the case, we fitted a Gaussian distribution to data that was sampled from a Mixture of Gaussians (Fig. 5a, left). The optimal C2ST between these two distributions is 0.65 (which can be computed because Gaussians and Mixtures of Gaussians allow evaluating densities). If the C2ST is estimated with a neural network, however, we observe that this C2ST can be systematically underestimated: for example, when only few samples from the data and generative model are available, the neural network predicts a C2ST of 0.5 (Fig. 5a, top right)—in other words, it predicts that the generative model and the data follow the same distribution. Similarly, if the neural network is not expressive enough, e.g., with too few hidden units, the classifier will also return a low C2ST, around 0.5 (Fig. 5a, bottom right). These issues can make the C2ST easy to misuse. In many cases, reporting a low C2ST is desirable for generative models since it indicates that the model perfectly matches data, but one can achieve a low C2ST simply by not investing sufficient time into obtaining a strong classifier.
@@
We previously argued that the C2ST is an interpretable measure—while this is generally true, the C2ST can sometimes be surprisingly high even if the generative model seems well aligned with the data. For example, when the generative model aligns very well with the data for every marginal, the C2ST can still be high if the data is high-dimensional (Fig. 5b). Because of this, it can be difficult to achieve low C2ST values on high-dimensional data. To further demonstrate this, we fitted a Gaussian distribution and a mixture of 20 Gaussian distributions to the ‘ones’ of the MNIST dataset. Although the Mixture of Gaussians (Fig. 5c, bottom row) looks better than a single Gaussian (Fig. 5c, middle row), both densities have a C2ST of 1.0 to the data (obtained with a ResNet on $\approx$4k held-out test datapoints).
@@
While we focus on a standard C2ST definition by using classification accuracy as the C2ST distance (Lopez-Paz & Oquab, 2017), any other performance measure for binary classification could be used (Raschka, 2014). Kim et al. (2019) argues that classic accuracy is sub-optimal due to the “binarization” of the class probabilities and proceeds to instead use the mean squared error between the predicted and “target” value of 0.5. Other approaches instead construct a likelihood ratio statistic (Pandeva et al., 2024). Additionally, instead of using the estimated class probabilities, Cheng & Cloninger (2022) consider using the average difference in logits (i.e., activations in the last hidden layer).
We note that the learned classifier in C2ST can be applied to estimations of a density ratio $\frac{p(x)}{q(x)}$, also known as the likelihood ratio trick (Hastie et al., 2001, Sugiyama et al., 2012). Density ratio estimation has attracted a great deal of attention in the statistics and machine learning communities since it can be employed for estimating divergences between two distributions, such as the Kullback–Leibler divergence (Titsias & Ruiz, 2019, Huszár, 2017) and Pearson divergence (Srivastava et al., 2020).
@@ 2.3 Kernel-based: maximum mean discrepancy (MMD)
MMD is a popular distance metric that is applicable to a variety of data domains, including high-dimensional continuous data spaces, strings of text, and graphs (Borgwardt et al., 2006, Gretton et al., 2012a, Muandet et al., 2017). It has been used to evaluate generative models (Sutherland et al., 2021, Borji, 2019, Lueckmann et al., 2021) and also has the ability to indicate where the model and the true distribution differ (Lloyd & Ghahramani, 2015). The distance provided by MMD can straightforwardly be used to test whether the difference between two sets of high-dimensional samples is statistically significant (Gretton et al., 2012a).
To assess whether two sets of samples are drawn from the same distribution, MMD makes use of a kernel function to (implicitly) embed the samples via an embedding function $\phi$, also called a feature map. If we choose the right kernel, we can end up embedding our samples in a space where the properties of the underlying distributions are easily compared. We will motivate the use of the kernel in MMD by illustrating different explicit embeddings before introducing the implicit embedding via a kernel $k$. Note that this explanation is inspired by Sutherland (2019).
@@
Figure 6: Maximum mean discrepancy (MMD). (a) Two example distributions $p_{1}(x)$, $p_{2}(x)$ and observed data $p_{true}(x)$ that we want to compare. (b) MMD can be defined as the difference between the expectations of some embedding function $\phi(x)$. If we take the identity as embedding ($\phi^{(1)}(x)=x$; left), we end up computing the differences between the means of the distributions, which are all equal for the three distributions. If we add a quadratic feature ($\phi^{(2)}(x)=[x,x^{2}]^{\mathsf{T}}$; right), we can distinguish distributions with different variances. Note that we still have $\mathsf{MMD}^{2}[\phi^{(2)},p_{2},p_{true}]=0$, despite $p_{2}$ being different from $p_{true}$ (c) Using the kernel trick we can avoid computing the embeddings all together but use implicit embeddings that capture all relevant features of the distributions.
If we have samples of real numbers from two distributions $p_{1}$ and $p_{2}$ (Fig. 6a), there are various embedding functions $\phi$ we could use to compare these. The simplest possible function $\phi^{(1)}:\mathbb{R}\rightarrow\mathbb{R}$ is the identity mapping $\phi^{(1)}(x)=x$ (Fig 6b, left). However, in this case, the MMD will simply be the absolute difference between the means (first moments) of the distributions (for details, see Section A.2):
@@
This does not yet allow us to discriminate different distributions with equal means (all distributions in Fig. 6). If we now expand our embedding with a quadratic term, $\phi^{(2)}:\mathbb{R}\rightarrow\mathbb{R}^{2}$ as $\phi^{(2)}(x)=\begin{bmatrix}x\\ x^{2}\end{bmatrix}$ (Fig 6b, right), the MMD yields (for details, see Section A.2)
@@
We can now rewrite the inner product $\langle\phi(x),\phi(x^{\prime})\rangle$ in terms of a kernel function $k$: $\langle\phi(x),\phi(x^{\prime})\rangle=k(x,x^{\prime})$. Thus, if we can find a kernel for our feature map, we can avoid explicitly computing the features altogether and instead directly compute
@@
Evaluating the kernel function instead of explicitly calculating the features is often called the kernel trick (Fig. 6c). If we can define a kernel whose corresponding embedding captures all, potentially infinitely many moments, we would have an MMD that is zero only if two distributions are exactly equal and the MMD becomes a metric. These kernels are called characteristic (Section A.2Gretton et al. 2012a), and include the commonly used Gaussian kernel: $k_{G}(x,x^{\prime})=\exp(-\frac{\lVert x-x^{\prime}\rVert^{2}}{2\sigma^{2}})$. A number of other kernels can also be used (see, e.g., Sriperumbudur et al. 2009). For instance, using a kernel induced by the Euclidean distance, MMD can be shown to be equivalent to the standard energy distance (Székely & Rizzo, 2013). In fact, a wider equivalence between MMD and the generalized energy distance has been established, using distance-induced kernels (Sejdinovic et al., 2013).
@@
Typically, the kernel version of MMD is used, which is straightforwardly estimated with its empirical, unbiased, estimate:
@@
MMD in this form can be applied to many forms of data as long as we can define a kernel, which can include graphs (Vishwanathan et al., 2010, Gärtner, 2003) or strings of text (Lodhi et al., 2002), in addition to vectors and matrices.
When we estimate the MMD with a finite number of samples, the selection of the right kernel and its parameters becomes crucial. For example, when using a Gaussian kernel, one has to choose the bandwidth $\sigma$. The MMD approaches zero if we take $\sigma$ to be close to zero (then $k_{G}(x,x^{\prime})=1\;\mathsf{if}\;x=x^{\prime}\;\mathsf{else}\;k_{G}(x,x^{\prime})\rightarrow 0$) or if $\sigma$ is large (then $k_{G}(x,x^{\prime})\rightarrow 1\;\forall\;x,x^{\prime}$; Gretton et al. 2012a). A common heuristic to remedy this parameter choice is picking the bandwidth based on the scale of the data. The median heuristic sets the bandwidth to the median distance between points in the aggregate sample (Gretton et al., 2012a). Another common approach is based on cross-validation, or data splitting (Gretton et al., 2012a; b, Jitkrittum et al., 2016, Sutherland et al., 2021): The dataset is divided, with a hold-out set used for kernel selection, and the other part used for evaluating MMD. While the data splitting method does not involve any heuristic, it can lead to errors in MMD since it reduces the number of data points available for estimating the MMD. Recent work attempts to choose hyperparameters without employing data splitting (Biggs et al., 2023, Schrab et al., 2023, Kübler et al., 2022b; a).
We often aim for a kernel that captures the (dis)similarity between the data points well, such a kernel can be domain-specific or specifically designed for downstream analysis tasks. The similarity between two strings (e.g., DNA sequences or text) can, for instance, be estimated by looking at the frequency of small subsequences (Leslie et al., 2001, Lodhi et al., 2002). Furthermore, it is possible to aggregate simpler kernels into a more expressive one (Gretton et al., 2012b), or to use a deep kernel (i.e., based on neural networks) that can exploit features of particular data modalities such as images (Liu et al., 2020, Gao et al., 2021).
@@ 2.4 Network-based: Embedding-space measures
Distribution comparisons on structured data spaces, such as a set of natural images, present unique challenges. Such data is usually high-dimensional (high-resolution images) and contains localized correlations. Furthermore, images of different object classes (such as airplanes and dogs) share low-level features in the form of edges and textural details but differ in semantic meaning. Similar challenges occur for time-series data, natural language text, and other complex data types (Smith & Smith, 2020, Jeha et al., 2021).
@@
Figure 7: Network-based metrics. Instead of directly computing distances in data space, complex data, e.g., natural images of dogs sampled from $p_{1}(x)$ and aircraft sampled from $p_{2}(y)$, are jointly embedded into a vector space. The embedding function can, for example, be a deep neural network. The resulting distributions in feature space are then compared by a classical measure of choice $D$.
@@
In this section, we rely on the example of natural images, but the presented framework generalizes to other data types. Naive distances would operate on a per-pixel basis, leading to scenarios where, for example, white dogs and black dogs are considered vastly different despite both being categorized as dogs. As we would like to have a distance measure that operates based on details relevant to the comparison, we can leverage neural networks trained on a large image dataset that captures features ranging from low-level to high-level semantic details: While earlier layers in a convolutional neural network focus on edge detection, color comparison, and texture detection, later layers learn to detect high-level features, such as a dog’s nose or the wing of an airplane, which thought to be relevant for a meaningful comparison. Embedding-based distances use these activations of neural network layers as an embedding to compare the image distributions. The most popular distance in this class is the Fréchet Inception Distance (FID; Heusel et al. 2017), used to evaluate generative models for images. The FID uses a convolutional neural network’s embeddings (specifically InceptionV3 (Szegedy et al., 2015)) to extract relevant features, applies a Gaussian approximation in the embedding space, and computes the Wasserstein distance on this approximation.
@@
A FID-like measure, in essence, requires a suitable embedding network $f:\mathcal{X}\rightarrow\mathbb{R}^{d}$, where $f$ transforms the data from the original high-dimensional space $\mathcal{X}$ into a lower-dimensional, feature-rich representation in $\mathbb{R}^{d}$ (Fig. 7). Once the data samples are mapped into this reduced space through the embedding network, the two sets of embedded samples can be compared using the appropriate distance. When evaluating generative models for natural images, it is common to approximate the embedded distributions with Gaussian distributions by estimating their respective mean $\mu$ and covariances $\Sigma$. Under this Gaussian approximation, the squared Wasserstein distance (also known as the Fréchet distance) can be analytically computed as
$W^{2}((\mu_{1},\Sigma_{1}),(\mu_{2},\Sigma_{2}))=\|\mu_{1}-\mu_{2}\|^{2}+\text{Tr}\left(\Sigma_{1}+\Sigma_{2}-2\left(\Sigma_{1}\Sigma_{2}\right)^{\frac{1}{2}}\right).$ (4)
@@
In principle, any appropriate metric can be used in place of the Fréchet distance. For instance, Jayasumana et al. (2023), Bińkowski et al. (2018), Xu et al. (2018) use MMD as a metric in the embedding space, and the MMD-based Inception Distance is often referred to as Kernel Inception Distance (KID). KID is known to have some advantages over FID: unlike FID, KID has a simple, unbiased estimator with better sample efficiency (Jayasumana et al., 2023), and does not assume any parametric forms for the distributions. Since KID involves MMD, we must carefully select the proper kernel and its hyperparameter when applying it. A related and commonly used quality measure for images is the Inception Score (Salimans et al., 2016). In contrast to the FID, this measure uses the average InceptionV3 predicted class probabilities and compares them with the true marginal class distribution. Note that while both this score and the FID can agree with traditional distances (e.g., certain divergences), they might evaluate models differently (Betzalel et al., 2022); see Barratt & Sharma (2018) for further limitations of the Inception Score.
@@
One of the biggest limitations is the requirement of a suitable embedding network. Newer and more robust networks, such as the image network of the CLIP (Radford et al., 2021) vision-language model, provide better and more semantically consistent embeddings (Betzalel et al., 2022, Jayasumana et al., 2023) than the InceptionV3 network. However, as the embedding network is generally non-injective, identical distributions in the embedding space may not necessarily translate to identical distributions in the original space. Previous research has demonstrated the FID’s sensitivity to preprocessing such as image resizing and compression (Parmar et al., 2022). Additionally, FID estimates are biased for finite sample sizes, making comparisons unreliable due to dependency on the generative model. However, methods to obtain a more unbiased estimate have been proposed ($\mathsf{FID}_{\infty}$; Chong & Forsyth 2020, Betzalel et al. 2022).
3 Comparing distances: sample size, dimensionality, and more